OpenStudy (anonymous):

Solve the system using the substitution method. 3x – 4y =10 2y = 2x – 6

OpenStudy (anonymous):

so far, I lined it up so the y's and x's are on top of each other. -4y+3x=10 2y-2x= 6

OpenStudy (anonymous):

I always have trouble deciding which variable to solve for first.

OpenStudy (anonymous):

use bottom one to solve for y

OpenStudy (anonymous):

can you write it out for me

OpenStudy (anonymous):

y=(2x-6)/2 divide both sides by 2. Perentheses needed to indicate you dividing the whole thing. (2x-6)/2 not same as 2x-6/2. y=x-3 then finally you can divide the 2 into 2 and -6

OpenStudy (anonymous):

which way are u doing it? substitution, elimination?

OpenStudy (anonymous):

your q says to use substitution :)

OpenStudy (anonymous):

wait why did you write y= (2x-6)/2 ?

OpenStudy (anonymous):

solved for the bottom equation thats why

OpenStudy (anonymous):

oh sorry woops

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

REPOST ________ y=(2x-6)/2 divide both sides by 2. Perentheses needed to indicate you dividing the whole thing. (2x-6)/2 not same as 2x-6/2. y=x-3 then finally you can divide the 2 into 2 and -6

OpenStudy (anonymous):

you didn't really show how u go it though..

OpenStudy (anonymous):

got*

OpenStudy (anonymous):

how do you go from y=(2x-6)/2 to y= x -3

OpenStudy (anonymous):